# ARECHO: Autoregressive Evaluation via Chain-Based Hypothesis Optimization for Speech Multi-Metric Estimation

### v1

@@ F.2 Metrics Information
The metrics information is detailed in Table 456, and 7.
@@
    • 4 metrics have unbounded ranges (,)($\infty$,$\infty$)( - ∞ , ∞ ) (e.g., SE-SI-SNR, SE-SDR).
    • 3 metrics are semi-bounded in $[0,\infty)$ (e.g., SRMR, SWR/SCR, Predicted Text Length).
@@
No. Name Type Range Model Based Target Direction Reference
@@
    • 3 metrics span $[-1,1]$ (e.g., F0-CORR, D-BERT, NCM).
    • Several metrics are unbounded or semi-bounded in $[0,\infty)$ or (,)($\infty$,$\infty$)( - ∞ , ∞ ), particularly SNR- and SDR-based metrics.
@@
No. Name Type Range Model Based Target Direction Reference
@@ 23 Normalized-Covariance Measure (NCM) numerical [-1, 1] $\uparrow$ Chen & Loizou (2010)
    • 2 metrics range from $[-1,1]$ (EMO-SIM, SPK-SIM), reflecting cosine similarity scales.
    • 2 metrics are semi-unbounded in $[0,\infty)$ (WER, CER).
@@
No. Name Type Range Model Based Target Direction Reference
@@ 7 Log Likelihood Ratio (LLR) numerical [0, inf) $\uparrow$ Hu & Loizou (2007a)
    • RT60 and Reference Text Length are semi-bounded in $[0,\infty)$.
    • SNR Simulation is unbounded in (,)($\infty$,$\infty$)( - ∞ , ∞ ), representing real-world variability in noise conditions.
@@
No. Name Type Range Target Direction Reference
@@ Appendix H Complete Main Experimental Table
Base Dev. UniVERSA 160.06 5.17 4.13 0.69 0.34 0.27 0.68 0.43 0.47 0.42
UniVERSA-T 40.95 2.70 1.62 0.78 0.41 0.34 0.70 0.49 0.50 0.46
ARECHO 25.73 2.16 1.27 0.86 0.43 0.36 0.71 0.52 0.53 0.51
Enhanced UniVERSA 61.54 4.22 3.48 0.71 0.36 0.27 0.69 0.43 0.48 0.43
UniVERSA-T 27.34 2.65 1.60 0.81 0.42 0.34 0.70 0.49 0.51 0.47
ARECHO 20.58 2.09 1.32 0.84 0.42 0.35 0.72 0.52 0.54 0.51
Corrupted UniVERSA 170.65 4.84 3.74 0.61 0.32 0.24 0.70 0.47 0.50 0.46
UniVERSA-T 77.72 2.91 1.70 0.74 0.41 0.34 0.71 0.52 0.53 0.50
ARECHO 44.22 2.37 1.29 0.82 0.42 0.35 0.72 0.56 0.56 0.55
Synthesized UniVERSA 58.79 3.82 3.29 0.76 0.37 0.27 0.69 0.45 0.49 0.45
UniVERSA-T 8.10 1.52 0.91 0.84 0.42 0.34 0.72 0.52 0.53 0.50
ARECHO 4.99 1.13 0.58 0.91 0.45 0.39 0.79 0.67 0.66 0.65
Avg. Test UniVERSA 96.99 4.29 3.50 0.69 0.35 0.26 0.69 0.45 0.49 0.45
UniVERSA-T 37.72 2.36 1.40 0.79 0.41 0.34 0.71 0.51 0.52 0.49
ARECHO 23.26 1.86 1.06 0.86 0.43 0.36 0.74 0.58 0.59 0.57
Scale Dev. UniVERSA 116.01 3.54 2.28 0.89 0.45 0.37 0.73 0.52 0.52 0.49
UniVERSA-T 27.98 2.39 1.22 0.86 0.43 0.37 0.74 0.54 0.54 0.52
ARECHO 29.61 2.49 1.32 0.85 0.43 0.37 0.75 0.54 0.54 0.52
Enhanced UniVERSA 43.05 2.53 1.81 0.84 0.42 0.34 0.72 0.49 0.51 0.47
UniVERSA-T 69.94 3.99 2.17 0.80 0.43 0.36 0.74 0.53 0.53 0.50
ARECHO 32.63 2.86 1.53 0.83 0.43 0.37 0.75 0.56 0.55 0.53
Corrupted UniVERSA 151.97 3.69 2.26 0.88 0.44 0.37 0.75 0.57 0.57 0.54
UniVERSA-T 39.80 2.42 1.18 0.77 0.43 0.37 0.76 0.58 0.57 0.54
ARECHO 34.37 2.32 1.10 0.84 0.43 0.38 0.77 0.59 0.58 0.56
Synthesized UniVERSA 6.46 1.49 1.00 0.84 0.41 0.32 0.71 0.48 0.51 0.47
UniVERSA-T 8.23 1.49 0.94 0.84 0.42 0.34 0.73 0.50 0.52 0.49
ARECHO 8.63 1.49 0.90 0.85 0.43 0.36 0.75 0.56 0.55 0.54
Avg. Test UniVERSA 67.16 2.57 1.69 0.85 0.43 0.35 0.73 0.51 0.53 0.50
UniVERSA-T 39.32 2.63 1.43 0.80 0.43 0.36 0.74 0.53 0.54 0.51
ARECHO 25.21 2.22 1.18 0.84 0.43 0.37 0.76 0.57 0.56 0.54
To provide a more comprehensive assessment of model performance, we expand our evaluation with additional regression metrics (i.e., Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Spearman’s Rank Correlation Coefficient (SRCC)) as well as classification metrics (i.e., Precision and Recall) complementing the original metrics presented in Table 9.
@@
Appendix I Ablation Experiments
@@ I.1 The Effect of Token Size
Table 10: Ablation study: the effect of token size for ARECHO.
Domain Model Beam Size Regression Metrics Classification Metrics
Dev. UniVERSA-T 500 27.98 2.39 1.22 0.86 0.43 0.37 0.74 0.54 0.54 0.52
1,000 50.46 3.21 1.63 0.82 0.42 0.36 0.73 0.53 0.52 0.50
ARECHO 500 29.61 2.49 1.32 0.85 0.43 0.37 0.75 0.54 0.54 0.52
1,000 30.86 2.53 1.33 0.84 0.43 0.37 0.74 0.55 0.54 0.51
Enhanced UniVERSA-T 500 69.94 3.99 2.17 0.80 0.43 0.36 0.74 0.53 0.53 0.50
1,000 59.33 3.80 2.07 0.81 0.42 0.35 0.73 0.53 0.53 0.50
ARECHO 500 32.63 2.86 1.53 0.83 0.43 0.37 0.75 0.56 0.55 0.53
1,000 43.32 3.17 1.70 0.81 0.43 0.36 0.75 0.54 0.54 0.52
Corrupted UniVERSA-T 500 39.80 2.42 1.18 0.77 0.43 0.37 0.76 0.58 0.57 0.54
1,000 60.73 2.66 1.29 0.82 0.43 0.37 0.75 0.57 0.56 0.54
ARECHO 500 34.37 2.32 1.10 0.84 0.43 0.38 0.77 0.59 0.58 0.56
1,000 31.68 2.02 1.07 0.84 0.43 0.37 0.76 0.59 0.58 0.56
Synthesized UniVERSA-T 500 8.23 1.49 0.94 0.84 0.42 0.34 0.73 0.50 0.52 0.49
1,000 8.50 1.53 0.92 0.83 0.41 0.34 0.73 0.51 0.52 0.49
ARECHO 500 8.63 1.49 0.90 0.85 0.43 0.36 0.75 0.56 0.55 0.54
1,000 7.31 1.41 0.83 0.85 0.43 0.36 0.75 0.57 0.55 0.53
@@ I.2 The Effect of Beam Search
The Synthesized domain, which represents our most controlled experimental setting, shows the best overall performance across all metrics. Here, beam size 2 achieves the optimal results with the lowest MSE (4.68) and RMSE (1.11), while beam size 3 yields the lowest MAE (0.57). The high correlation coefficients (LCC=0.91, SRCC=0.45, KTAU=0.39) across all beam sizes indicate that the model effectively captures the underlying relationships in this domain, regardless of the search strategy employed.
@@
Dev. 1 25.73 2.16 1.27 0.86 0.43 0.36 0.71 0.52 0.53 0.51
2 25.12 2.14 1.27 0.86 0.43 0.36 0.71 0.52 0.53 0.50
3 25.52 2.13 1.26 0.86 0.43 0.36 0.71 0.52 0.53 0.50
4 26.07 2.22 1.28 0.85 0.43 0.36 0.71 0.52 0.53 0.50
Enhanced 1 20.58 2.09 1.32 0.84 0.42 0.35 0.72 0.52 0.54 0.51
2 22.67 2.19 1.36 0.84 0.43 0.35 0.72 0.52 0.54 0.52
3 22.98 2.20 1.37 0.84 0.43 0.35 0.72 0.52 0.54 0.51
4 22.63 2.18 1.36 0.84 0.43 0.35 0.72 0.52 0.54 0.52
Corrupted 1 44.22 2.37 1.29 0.82 0.42 0.35 0.72 0.56 0.56 0.55
2 41.02 2.26 1.23 0.83 0.42 0.35 0.73 0.56 0.56 0.55
3 40.91 2.26 1.23 0.83 0.42 0.35 0.73 0.56 0.56 0.55
4 41.10 2.26 1.23 0.83 0.42 0.35 0.73 0.56 0.56 0.55
Synthesized 1 4.99 1.13 0.58 0.91 0.45 0.39 0.79 0.67 0.66 0.65
2 4.68 1.11 0.58 0.91 0.45 0.39 0.79 0.67 0.67 0.65
3 4.85 1.12 0.57 0.91 0.45 0.39 0.79 0.67 0.66 0.64
4 4.95 1.13 0.58 0.91 0.45 0.39 0.79 0.67 0.66 0.64
Appendix J Dependency Analysis - Dynamic Dependency Analysis

### v2

@@ F.2 Metrics Information
The metrics information is detailed in Table 567, and 8.
@@
    • 4 metrics have unbounded ranges $(-\infty,\infty)$ (e.g., SE-SI-SNR, SE-SDR).
    • 3 metrics are semi-bounded in $0,\infty)$ (e.g., SRMR, SWR/SCR, Predicted Text Length).
@@
No. Name Type Range
Model
Based
Target
Direction
Reference
@@
    • 3 metrics span [1,1][-1,1] (e.g., F0-CORR, D-BERT, NCM).
    • Several metrics are unbounded or semi-bounded in $0,\infty)$ or $(-\infty,\infty)$, particularly SNR- and SDR-based metrics.
@@
No. Name Type Range
Model
Based
Target
Direction
Reference
@@ 23 Normalized-Covariance Measure (NCM) numerical [-1, 1] $\uparrow$ Chen & Loizou (2010)
    • 2 metrics range from [1,1][-1,1] (EMO-SIM, SPK-SIM), reflecting cosine similarity scales.
    • 2 metrics are semi-unbounded in $0,\infty)$ (WER, CER).
@@
No. Name Type Range
Model
Based
Target
Direction
Reference
@@ 7 Log Likelihood Ratio (LLR) numerical [0, inf) $\uparrow$ Hu & Loizou (2007a)
    • RT60 and Reference Text Length are semi-bounded in $0,\infty)$.
    • SNR Simulation is unbounded in $(-\infty,\infty)$, representing real-world variability in noise conditions.
@@
No. Name Type Range
Target
Direction
Reference
@@ Appendix H Complete Main Experimental Table
MSE ($\downarrow$) RMSE ($\downarrow$) MAE ($\downarrow$) BMAE ($\downarrow$) LCC ($\uparrow$) SRCC ($\uparrow$) KTAU ($\uparrow$) Acc ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$) F1 ($\uparrow$)
Base Dev. UniVERSA 160.06 5.17 4.13 5.08 0.69 0.68 0.53 0.68 0.43 0.47 0.42
UniVERSA-T 40.95 2.70 1.62 1.96 0.78 0.82 0.68 0.70 0.49 0.50 0.46
ARECHO 25.73 2.16 1.27 1.51 0.86 0.86 0.72 0.71 0.52 0.53 0.51
Enhanced UniVERSA 61.54 4.22 3.48 3.61 0.71 0.71 0.54 0.69 0.43 0.48 0.43
UniVERSA-T 27.34 2.65 1.60 1.75 0.81 0.84 0.68 0.70 0.49 0.51 0.47
ARECHO 20.58 2.09 1.32 1.43 0.84 0.85 0.69 0.72 0.52 0.54 0.51
Corrupted UniVERSA 170.65 4.84 3.74 4.79 0.61 0.63 0.48 0.70 0.47 0.50 0.46
UniVERSA-T 77.72 2.91 1.70 2.02 0.74 0.81 0.67 0.71 0.52 0.53 0.50
ARECHO 44.22 2.37 1.29 1.52 0.82 0.84 0.70 0.72 0.56 0.56 0.55
Synthesized UniVERSA 58.79 3.82 3.29 3.36 0.76 0.73 0.54 0.69 0.45 0.49 0.45
UniVERSA-T 8.10 1.52 0.91 0.97 0.84 0.83 0.68 0.72 0.52 0.53 0.50
ARECHO 4.99 1.13 0.58 0.61 0.91 0.91 0.78 0.79 0.67 0.66 0.65
Avg. Test UniVERSA 96.99 4.29 3.50 4.21 0.69 0.69 0.52 0.69 0.45 0.49 0.45
UniVERSA-T 37.72 2.36 1.40 1.68 0.79 0.83 0.68 0.71 0.51 0.52 0.49
ARECHO 23.26 1.86 1.06 1.27 0.86 0.87 0.72 0.74 0.58 0.59 0.57
Scale Dev. UniVERSA 116.01 3.54 2.28 4.02 0.89 0.89 0.74 0.73 0.52 0.52 0.49
UniVERSA-T 27.98 2.39 1.22 1.43 0.86 0.86 0.75 0.74 0.54 0.54 0.52
ARECHO 29.61 2.49 1.32 1.55 0.86 0.87 0.76 0.75 0.54 0.54 0.52
Enhanced UniVERSA 43.05 2.53 1.81 1.93 0.84 0.84 0.67 0.72 0.49 0.51 0.47
UniVERSA-T 69.94 3.99 2.17 2.33 0.80 0.86 0.71 0.74 0.53 0.53 0.50
ARECHO 32.63 2.86 1.53 1.67 0.83 0.87 0.73 0.75 0.56 0.55 0.53
Corrupted UniVERSA 151.97 3.69 2.26 4.11 0.88 0.89 0.75 0.75 0.57 0.57 0.54
UniVERSA-T 39.80 2.42 1.18 1.37 0.77 0.86 0.74 0.76 0.58 0.57 0.54
ARECHO 34.37 2.32 1.10 1.35 0.84 0.87 0.76 0.77 0.59 0.58 0.56
Synthesized UniVERSA 6.46 1.49 1.00 1.05 0.84 0.82 0.65 0.71 0.48 0.51 0.47
UniVERSA-T 8.23 1.49 0.94 0.99 0.84 0.83 0.68 0.73 0.50 0.52 0.49
ARECHO 8.63 1.49 0.90 0.94 0.85 0.85 0.72 0.75 0.56 0.55 0.54
Avg. Test UniVERSA 67.16 2.57 1.69 2.78 0.86 0.86 0.70 0.73 0.51 0.53 0.50
UniVERSA-T 39.32 2.63 1.43 1.53 0.82 0.86 0.72 0.74 0.53 0.54 0.51
ARECHO 25.21 2.22 1.18 1.38 0.85 0.87 0.74 0.76 0.57 0.56 0.54
To provide a more comprehensive assessment of model performance, we expand our evaluation with additional regression metrics (i.e., Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), Balanced Mean Absolute Error (BMAE) in Baccianella et al. (2009) and Spearman’s Rank Correlation Coefficient (SRCC)) as well as classification metrics (i.e., Precision and Recall) complementing the original metrics presented in Table 10.
@@
Appendix I Ablation Experiments and Model Analysis
Table 11: Ablation study: the effect of token size for ARECHO.
Domain Model Token Bins Regression Metrics Classification Metrics
Dev. UniVERSA-T 500 27.98 2.39 1.22 0.86 0.86 0.75 0.74 0.54 0.54 0.52
1,000 50.46 3.21 1.63 0.82 0.84 0.72 0.73 0.53 0.52 0.50
ARECHO 500 29.61 2.49 1.32 0.85 0.86 0.76 0.75 0.54 0.54 0.52
1,000 30.86 2.53 1.33 0.84 0.85 0.74 0.74 0.55 0.54 0.51
Enhanced UniVERSA-T 500 69.94 3.99 2.17 0.80 0.86 0.71 0.74 0.53 0.53 0.50
1,000 59.33 3.80 2.07 0.81 0.85 0.70 0.73 0.53 0.53 0.50
ARECHO 500 32.63 2.86 1.53 0.83 0.87 0.73 0.75 0.56 0.55 0.53
1,000 43.32 3.17 1.70 0.81 0.85 0.72 0.75 0.54 0.54 0.52
Corrupted UniVERSA-T 500 39.80 2.42 1.18 0.77 0.86 0.74 0.76 0.58 0.57 0.54
1,000 60.73 2.66 1.29 0.82 0.85 0.73 0.75 0.57 0.56 0.54
ARECHO 500 34.37 2.32 1.10 0.84 0.87 0.76 0.77 0.59 0.58 0.56
1,000 31.68 2.02 1.07 0.84 0.85 0.74 0.76 0.59 0.58 0.56
Synthesized UniVERSA-T 500 8.23 1.49 0.94 0.84 0.83 0.68 0.73 0.50 0.52 0.49
1,000 8.50 1.53 0.92 0.83 0.83 0.65 0.73 0.51 0.52 0.49
ARECHO 500 8.63 1.49 0.90 0.85 0.85 0.72 0.75 0.56 0.55 0.54
1,000 7.31 1.41 0.83 0.85 0.85 0.71 0.75 0.57 0.55 0.53
@@ I.2 The Effect of Beam Search
Dev. 1 25.73 2.16 1.27 0.86 0.86 0.72 0.71 0.52 0.53 0.51
2 25.12 2.14 1.27 0.86 0.86 0.72 0.71 0.52 0.53 0.50
3 25.52 2.13 1.26 0.86 0.86 0.72 0.71 0.52 0.53 0.50
4 26.07 2.22 1.28 0.85 0.86 0.72 0.71 0.52 0.53 0.50
Enhanced 1 20.58 2.09 1.32 0.84 0.85 0.69 0.72 0.52 0.54 0.51
2 22.67 2.19 1.36 0.84 0.85 0.69 0.72 0.52 0.54 0.52
3 22.98 2.20 1.37 0.84 0.85 0.69 0.72 0.52 0.54 0.51
4 22.63 2.18 1.36 0.84 0.85 0.69 0.72 0.52 0.54 0.52
Corrupted 1 44.22 2.37 1.29 0.82 0.84 0.70 0.72 0.56 0.56 0.55
2 41.02 2.26 1.23 0.83 0.84 0.70 0.73 0.56 0.56 0.55
3 40.91 2.26 1.23 0.83 0.84 0.70 0.73 0.56 0.56 0.55
4 41.10 2.26 1.23 0.83 0.84 0.70 0.73 0.56 0.56 0.55
Synthesized 1 4.99 1.13 0.58 0.91 0.91 0.78 0.79 0.67 0.66 0.65
2 4.68 1.11 0.58 0.91 0.91 0.78 0.79 0.67 0.67 0.65
3 4.85 1.12 0.57 0.91 0.91 0.78 0.79 0.67 0.66 0.64
4 4.95 1.13 0.58 0.91 0.91 0.78 0.79 0.67 0.66 0.64
The Synthesized domain, which represents our most controlled experimental setting, shows the best overall performance across all metrics. Here, beam size 2 achieves the optimal results with the lowest MSE (4.68) and RMSE (1.11), while beam size 3 yields the lowest MAE (0.57). The high correlation coefficients (LCC=0.91) across all beam sizes indicate that the model effectively captures the underlying relationships in this domain, regardless of the search strategy employed.
Table 13: Teacher-forced decoding analysis. We compare the original baseline, autoregressive ARECHO, and teacher-forced decoding (which injects ground-truth values at each decoding step to suppress error propagation). We use (T.F.) to note model use teacher-forcing decoding.
Domain Model Regression Metrics Classification Metrics
MSE ($\downarrow$) RMSE ($\downarrow$) MAE ($\downarrow$) LCC ($\uparrow$) SRCC ($\uparrow$) KTAU ($\uparrow$) Acc ($\uparrow$) F1 ($\uparrow$)
Dev. ARECHO 40.95 2.16 1.27 0.86 0.86 0.72 0.71 0.51
ARECHO (T.F.) 25.08 1.80 0.94 0.88 0.89 0.77 0.71 0.50
Enhanced ARECHO 20.58 2.09 1.32 0.84 0.85 0.69 0.72 0.51
ARECHO (T.F.) 3.80 1.01 0.43 0.88 0.91 0.76 0.72 0.51
Corrupted ARECHO 44.22 2.37 1.29 0.82 0.84 0.70 0.72 0.55
ARECHO (T.F.) 8.58 1.22 0.56 0.83 0.85 0.70 0.72 0.55
Synthesized ARECHO 4.99 1.13 0.58 0.91 0.91 0.78 0.72 0.55
ARECHO (T.F.) 2.00 0.80 0.30 0.92 0.93 0.79 0.79 0.64
I.3 Error Propagation Analyais / Hybrid Decoding
To better understand the effect of error propagation in autoregressive decoding, we investigate the decoding with teacher-forcing, where ground-truth metric values are injected at each step during inference. This removes the accumulation of decoding errors and provides an upper bound on model performance under ideal conditioning.
As shown in Table 13, teacher-forced decoding consistently improves over autoregressive ARECHO, yielding a 25–50% reduction in prediction error. This quantifies the impact of error propagation in sequential metric estimation. The effect is most pronounced in the corrupted and synthesized conditions, where domain shifts amplify early mistakes.
These findings highlight a practical trade-off: while teacher-forcing suppresses propagation and approximates an upper bound, it relies on ground-truth access and thus forfeits the fully reference-free nature of ARECHO.
I.4 Analysis on MOS Prediction
To further analyze the per-metric performance, we conduct a detailed comparison to investigate how inter-metric dependencies affect the modeling of MOS scores. As shown below, ARECHO consistently outperforms UniVERSA-T across multiple datasets, highlighting the benefit of inter-metric prediction and autoregressive dependency modeling.
Table 14: Comparison on MOS-related metrics across four public datasets. ARECHO shows consistent improvements in both regression and rank correlation measures over UniVERSA-T.
Metric Model Regression Metrics Correlation Metrics
MSE ($\downarrow$) RMSE ($\downarrow$) MAE ($\downarrow$) LCC ($\uparrow$) SRCC ($\uparrow$) KTAU ($\uparrow$)
NISQA-MOS UniVERSA-T 0.41 0.64 0.43 0.86 0.83 0.69
ARECHO 0.30 0.55 0.37 0.89 0.88 0.72
UTMOS UniVERSA-T 0.06 0.25 0.19 0.97 0.97 0.85
ARECHO 0.05 0.22 0.16 0.97 0.97 0.87
PLCMOS UniVERSA-T 0.37 0.61 0.41 0.86 0.86 0.69
ARECHO 0.32 0.56 0.38 0.89 0.89 0.72
URGENT-MOS UniVERSA-T 0.20 0.45 0.35 0.78 0.78 0.62
ARECHO 0.26 0.51 0.40 0.74 0.75 0.61
Overall, ARECHO demonstrates notable improvements in MSE and MAE, particularly for NISQA-MOS, UTMOS, and PLCMOS. The correlation gains (LCC/SRCC/KTAU) further suggest that ARECHO effectively captures cross-metric dependencies that are beneficial for reliable MOS prediction.
Table 15: DNSMOS ablation on ARECHO. “Full” denotes multi-metric training, while “Only” refers to single-metric prediction.
Metric Model Regression Metrics Correlation Metrics
MSE ($\downarrow$) RMSE ($\downarrow$) MAE ($\downarrow$) LCC ($\uparrow$) SRCC ($\uparrow$) KTAU ($\uparrow$)
DNSMOS ARECHO-full 0.04 0.21 0.16 0.85 0.81 0.65
ARECHO-only 0.04 0.20 0.15 0.86 0.83 0.66
Interestingly, as shown in Table 15, the difference between ARECHO-full (multi-metric prediction) and ARECHO-only (single-metric prediction) is marginal for DNSMOS. This suggests that the benefit of inter-metric conditioning varies by dataset complexity and metric interdependence: tasks with stronger perceptual correlation (e.g., MOS-related metrics) benefit more from autoregressive chaining, whereas narrowly defined metrics like DNSMOS can be learned independently.
Appendix J Task-Oriented Metric Subset Analysis
To further examine whether training with all metrics is always optimal and whether unrelated metrics might negatively affect task-relevant ones, we conduct targeted experiments comparing subset and fullset training configurations across two representative domains: speech synthesis and speech enhancement. The results show that ARECHO remains robust under both settings, with diverse metric supervision generally improving or maintaining performance.
The dynamic classifier chain design enables each metric to be predicted at a random position within the autoregressive chain, conditioned on a variable subset of preceding metrics. This stochastic ordering allows ARECHO to capture informative inter-metric dependencies while factorizing away irrelevant information, preventing negative transfer even under mixed supervision.
All experiments (UniVERSA baseline, Subset, and Fullset) are trained on the same Base dataset described in Sec. 4.1 and detailed in Appendix E. The only difference lies in the set of target metrics:
  • Speech Synthesis Subset: UTMOS, F0 correlation, WER, Language, etc.
  • Speech Enhancement Subset: DNSMOS, PESQ, RIR Room Size, etc.
  • Fullset: Union of all metrics from both domains.
No additional data is introduced for Fullset training; any performance differences therefore reflect the effect of training objectives rather than data volume.
Not all utterances contain all metrics in the base dataset due to data-source variability. For example:
  • Only 13.22% of samples contain NISQA Real MOS (human-annotated).
  • Nearly all have UTMOS and Language (non-intrusive model-based).
  • For enhancement data, $\sim$26.79% include intrusive metrics (PESQ/STOI/SDR), 80.44% have SRMR, and all have DNSMOS and Qwen-Recording Quality.
ARECHO naturally handles partial labels via its dynamic chain formulation, learning from whatever subset of metrics is available per sample. In contrast, UniVERSA relies on masking-based multi-task training, which reduces both data efficiency and inter-metric contextualization.
Table 16: Subset vs. fullset training comparison for speech synthesis metrics (Synthesized test set is used for evaluation).
Metric Setup MSE ($\downarrow$) RMSE ($\downarrow$) MAE ($\downarrow$)
NISQA Real MOS UniVERSA 1.22 1.10 0.85
Subset 0.43 0.65 0.41
Fullset 0.05 0.23 0.12
UTMOS UniVERSA 0.25 0.50 0.39
Subset 0.09 0.30 0.22
Fullset 0.04 0.20 0.13
Table 17: Subset vs. fullset training comparison for speech synthesis classification metrics (Synthesized test set is used for evaluation).
Metric Setup Acc. ($\uparrow$) F1 ($\uparrow$)
Language UniVERSA 0.90 0.88
Subset 0.96 0.96
Fullset 0.98 0.98
Table 18: Subset vs. fullset training comparison for speech enhancement metrics (Enhanced test set is used for evaluation).
Metric Setup MSE ($\downarrow$) RMSE ($\downarrow$) MAE ($\downarrow$)
SRMR UniVERSA 73.73 8.59 7.86
Subset 4.02 2.01 1.26
Fullset 1.83 1.35 0.96
DNSMOS UniVERSA 5.50 2.34 2.29
Subset 0.05 0.22 0.16
Fullset 0.05 0.22 0.16
STOI UniVERSA 0.16 0.41 0.39
Subset 0.04 0.06 0.03
Fullset 0.00 0.05 0.03
SDR UniVERSA 151.75 12.32 10.70
Subset 69.43 8.33 4.02
Fullset 19.83 4.45 2.93
PESQ UniVERSA 0.46 0.68 0.53
Subset 0.19 0.44 0.33
Fullset 0.17 0.41 0.31
Table 19: Subset vs. fullset training comparison for speech enhancement classification metrics (Enhanced test set is used for evaluation).
Metric Setup Acc. ($\uparrow$) F1 ($\uparrow$)
Qwen-Recording Quality UniVERSA 0.97 0.95
Subset 0.97 0.95
Fullset 0.97 0.95
The results shown in Table 16-19 confirm three main trends:
  • Fullset training outperforms Subset training for most metrics (NISQA MOS, SRMR, SDR) and performs on par for metrics such as DNSMOS and STOI.
  • No negative transfer is observed: unrelated metrics do not degrade the performance of task-relevant ones.
  • Dynamic classifier chain improves stability: randomizing dependency order during training allows decoupling of irrelevant signals while leveraging cross-metric correlations when beneficial.
Performance gains differ based on metric availability and difficulty:
Label Availability and Dependency Opportunity.
Metrics with limited supervision but strong inter-metric correlation (e.g., NISQA MOS) benefit the most from joint training. Densely available metrics (UTMOS, DNSMOS) show smaller improvements.
Metric Type and Modeling Difficulty.
Subjective and high-variance metrics gain more from multi-metric context, while objective or low-variance ones are largely self-sufficient.
Examples. NISQA MOS exhibits large improvements due to its sparse labels and correlation with perceptual quality scores. SDR benefits from complementary cues from PESQ, DNSMOS, and STOI, while UTMOS and DNSMOS remain stable due to high label coverage.
For further analysis of task-specific metric sensitivity, metrics can be grouped into:
  • Easy-dependent metrics: high correlation and availability (e.g., UTMOS, DNSMOS) that perform well without additional context.
  • Hard-dependent metrics: sparse or subjective metrics (e.g., NISQA MOS, STOI) that benefit strongly from auxiliary conditioning.
Overall, ARECHO demonstrates consistent performance across both homogeneous and heterogeneous metric sets. Joint training with diverse metric objectives does not harm, and often enhances, task-specific predictions. This stability stems from the dynamic classifier chain’s ability to balance inter-metric dependency and independence, enabling scalable and interpretable multi-domain evaluation.
Appendix K Dependency Analysis - Dynamic Dependency Analysis